# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

### v2

@@ 4 Architectures for Long-Horizon Vision-Language-Action Models
Context Extension. Given the limitations in context length of most LLMs, using off-the-shelf models is insufficient for processing long inputs such as those exceeding 1M tokens. We explore various long-context extension techniques that allows the model to generalize to longer input sequences without retraining from scratch. Specifically, we consider: Linear Interpolation(Chen et al., 2023): Rescales input positions to fit within the pretrained RoPE range by linearly interpolating positional indices; Dynamic Scaling(Chen et al., 2023): Adapts RoPE frequencies at runtime based on the input sequence length, using a linear rescaling to maintain consistent positional encoding behavior across varying lengths; YaRN(Peng et al., 2024): Dynamically interpolates attention frequencies during inference, balancing between pretrained and extrapolated positional regimes; LongRoPE(Ding et al., 2024): Augments RoPE with specially designed extrapolation functions, enabling robust generalization to long sequences without degrading attention quality. We apply these techniques during fine-tuning, at inference time, or both.
Context Parallelism. To further enhance the model’s ability to reason over long contexts, it is crucial to fine-tune on extended context inputs. However, the quadratic complexity of the attention mechanism makes it computationally infeasible to train directly over long sequences. To address this challenge, we employ Context Parallelism, a parallel training technique designed for efficient long-context modeling.
Context Parallelism leverages Ring Attention (Liu et al., 2023a), a novel parallel implementation of the attention layer. In Ring Attention, key-value (KV) shards are cyclically shuffled across devices, and partial attention scores are computed iteratively. This process is repeated until all KV shards have been incorporated on each device, ensuring complete attention coverage without the full memory cost of standard attention. By combining Context Parallelism with our dataset of extended long-context inputs, we are able to scale fine-tuning to substantially longer sequences, unlocking improved long-horizon reasoning capabilities.
@@ 5.1 Static Evaluation: Needle(s) in the Embodied Haystack
(a) Results of Needle in the Embodied Haystack (NiEH).
(b) Results of Needles in the Embodied Haystack (NsiEH).
Figure 4: Results of Needle(s) in the Embodied Haystack. The white dashed line denotes the maximum input context length of the model. Qwen2.5-VL was pre-trained initially with an 8K token context window and incrementally scaled up to 32K tokens in subsequent stages (Bai et al., 2025). The gray area indicates contexts not applicable (N/A) due to the model’s smaller image token size limiting sequences to under 128K tokens. Context Parallelism is applied to all experiments with the context size over 384K.
Results. Figure 4 presents the performance of LLaVA-OneVision 7B (Li et al., 2024a) model across various context extension methods. Linear Interpolation, Dynamic Scaling, and LongRoPE scaling all struggled with very long contexts beyond 128K tokens (the results of Linear Interpolation are excluded from Figure since it fails at all examples). YaRN consistently outperformed other methods, effectively handling contexts above 384K tokens, likely due to its architectural alignment with LLaVA-OneVision’s Qwen2 LM backbone, which employs RoPE and YaRN scaling during pretraining. YaRN performed best at moderate scaling factors (e.g., x4), however, further scaling to x8 and x16 did not yield additional gains. In particular, x16 slightly improved performance in the 256K–384K token range but led to degradation notably at shorter context sizes (¡64K), suggesting that excessive scaling may introduce instability and negatively impact performance. Overall, all methods fail beyond 512K tokens, highlighting a need for improved long-context methods.
Figure 4 also includes additional results from DeepSeek-VL 7B (Lu et al., 2024), Qwen2.5-VL 7B (Bai et al., 2025), and Gemini 2.0 Flash [^1] We used the gemini-2.0-flash-001 version for all experiments. (Google DeepMind, 2024). Each model processes images into tokens differently—DeepSeek (576 tokens/image), Qwen2.5 (121 tokens/image), and Gemini 2.0 Flash (258 tokens/image)—which consequently impacts maximum context lengths when transforming N(s)iEH sequences into tokenized inputs (DeepSeek-VL: 512K, Qwen2.5-VL: 128K, Gemini 2.0 Flash: 256K). DeepSeek-VL struggles beyond its 4K pretrained context limit. Qwen2.5-VL maintains stronger performance up to approximately 64K tokens, benefiting from its smaller per-image tokenization, but performance notably declines on NsiEH at longer contexts. Gemini 2.0 Flash performs robustly up to 8K tokens but degrades beyond 128K, especially on NsiEH tasks, highlighting room for improvement in complex long-range multimodal reasoning.
@@
Single vs Multi-evidence Reasoning. Comparing NiEH to the more challenging NsiEH task, we observe a significant performance drop in the multi-evidence setting. This is especially pronounced for mid-depth questions involving sparse or distant evidence (e.g., “Where was the Mug before you put it on the CounterTop?”) or questions requiring the aggregation of multiple clues (e.g., “How many times did you move the Apple?”), as shown in Figure 2(b). These results demonstrate that our NiEH and NsiEH tasks pose a substantial challenge to current long-context models and success requires both fine-grained temporal memory and multi-evidence reasoning across extended interactions.
5.2 Interactive Evaluation in $\infty$-Thor
To measure agent performance on our long-horizon test set, we conduct an interactive online evaluation using the AI2THOR for high-level planning tasks and the ManipulaTHOR for low-level manipulation tasks.
High-level Planning Task. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. We used 8 H100 GPUs with both tensor parallelism (Shoeybi et al., 2020) and pipeline parallelism (Huang et al., 2019) for the 32K context size, while Context Parallelism is employed for training on larger context sizes (64K and 130K). Additional training specifics are available in the Appendix. Figure 5 presents the accumulated rewards over time across six experimental configurations. The Planner trajectory serves as the performance upper bound.
Figure 5: Agent’s reward across different experimental configurations for high-level planning tasks. We compare (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, and (c) fine-tuning with different context lengths using Context Parallelism. (d) summarizes the most effective strategies, highlighting that exposure to longer contexts during training significantly improves performance. Non-planner models cannot generate valid actions after around 250 steps ($\approx$376K in context size). More configurations are in Figure 7 in Appendix.
In Figure 5(a), at a fixed scaling factor of x4, YaRN consistently achieves the highest performance, closely matching the Planner. Figure 5(b) shows that scaling YaRN further to x8 or x16 yields little additional gain. In Figure 5(c), fine-tuning with Context Parallelism extends context lengths to 64K and 130K tokens, enabling the model to learn much longer sequences (86 vs. 22 steps with 32K). Additional scaling at evaluation after fine-tuning (Figures 7(d,e) in Appendix) provides no improvement and may even degrade shorter-context performance. Overall, fine-tuning on long-trajectory datasets proves most effective (Figure 5(d)), while YaRN at x4 offers strong results when large-scale training data are unavailable.
Low-level Manipulation Task. Figure 6 shows agent rewards on low-level manipulation tasks using OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025). Ego-centric views are provided as image input, and the models generate the next action $(\Delta x,\Delta y,\Delta z)$ to control the robot arm. Both models perform below the Planner due to differences in robot arm configuration and out-of-distribution inputs, though SpatialVLA achieves slightly higher and more stable rewards than OpenVLA over long sequences. We provide more results and analysis in Appendix C.
@@ 6 Conclusion
We presented $\infty$-Thor, a new framework for long-horizon embodied tasks designed to advance long-context understanding in embodied AI. Our framework enables scalable synthesis of long, complex trajectories paired with high- and low-level action sequences, and supports both offline training and online interaction with the environment. As part of this framework, we introduced a novel embodied QA benchmark, Needle(s) in the Embodied Haystack, that challenges agents to reason over sparse, temporally distant visual evidence embedded within extended trajectories. To equip models for this setting, we explored architectural adaptations including interleaved Goal–State–Action modeling, context extension techniques such as YaRN and LongRoPE, and efficient fine-tuning via Context Parallelism. Our experiments demonstrate that exposure to longer contexts during training significantly improves model performance, and the limitation of existing context extension techniques struggle with long-context reasoning. We hope our framework and benchmark encourage further research into models capable of robust long-horizon reasoning under realistic, interactive environments.
Reproducibility statement
Experiments in this work were conducted using the publicly available simulators AI2-THOR and ManipulaTHOR, which allow for the deterministic replaying of all trajectories provided by $\infty$-Thor. We provide the complete source code, along with the generated datasets and configurations used for each experiment, and this will be made available in a public repository upon publication. Furthermore, all VLM models used in our experiments are open-source and available on Hugging Face. For the proprietary model Gemini Flash, we have declared the specific version of Gemini Flash used in the paper, and our experiments can be reproduced using the Google Cloud APIs. While the initial generation of trajectories involves some randomness in the sampling of target objects and destination positions, the resulting trajectories themselves are fully reproducible within the simulator environments. The detailed data processing steps and the experimental setup are further described in the Appendix.
@@ A.2 Constructing Long-Horizon Trajectories
To synthesize long-horizon trajectories, we construct each trajectory by sequentially chaining successful sub-tasks sampled from a predefined set of task templates. This process is illustrated in Algorithm 1. We begin by sampling a task template from a fixed task pool, which includes goal types such as pick and place simple, pick two obj and place, and pick and place with movable recep. Each sampled template requires relevant objects in the scene (e.g., pickupable items, target receptacles), which are then used to define the goal for that task.
@@
13:// Final sub-task with long-term object dependency
@@
Appendix B Training and Evaluation Details
B.1 Interactive Evaluation in $\infty$-Thor
Training. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. The model is trained using a next-action prediction objective, where only the action tokens are optimized, conditioned on the goal and state tokens. Table 5 summarizes the training specifications for different context lengths. For 32K training, we apply tensor parallelism with a degree of 4 and pipeline parallelism with a degree of 2, utilizing 8 H100 GPUs in total. Since pipeline parallelism requires the batch size to match the pipeline degree, we set the batch size to 2. For longer context lengths, we use context parallelism: 8-way for 64K (on 8 GPUs) and 16-way for 130K (on 16 GPUs). All models are fine-tuned for approximately 3 epochs with a learning rate of 1e-5, using the AdamW optimizer and a linear learning rate schedule with a 0.03 warmup ratio.
@@ B.1 Interactive Evaluation in $\infty$-Thor
Plan-Level Evaluation.
We evaluate agent performance using a plan-level framework, where each plan corresponds to a short sequence of actions aimed at achieving a specific intermediate sub-goal (e.g., navigating to an object, placing an item). A trajectory is composed of multiple such plans, executed sequentially. For the interactive evaluation, the agent is presented with the current plan’s goal along with the history of previous GT states and actions. Using this context, the agent predicts the next action and interacts step-by-step with the environment. The interaction continues until the current plan is either successfully completed or terminated due to failure (e.g., collisions or deadlocks). After each plan, the context is reset to include the GT actions and states from the completed portion of the trajectory, and the agent proceeds to the next plan. This ensures that each plan is evaluated independently, conditioned only on the correct prior history. The agent’s performance is measured via cumulative reward across all plans in the trajectory. Pseudocode for this evaluation procedure is provided in Algorithm 2.

### v3

@@ 4 Architectures and Training Considerations for Long-Horizon Embodied Reasoning
Context Extension and Context Parallelism. Given the limitations in context length of most VLMs, using off-the-shelf models is insufficient for processing long inputs such as those exceeding 1M tokens, particularly in our interleaved Goal-State-Action setting. We explore various long-context extension techniques that allows the model to generalize to longer input sequences without retraining from scratch. Specifically, we consider: Linear Interpolation (Chen et al., 2023): Rescales input positions to fit within the pretrained RoPE range by linearly interpolating positional indices; Dynamic Scaling (Chen et al., 2023): Adapts RoPE frequencies at runtime based on the input sequence length, using a linear rescaling to maintain consistent positional encoding behavior across varying lengths; YaRN (Peng et al., 2024): Dynamically interpolates attention frequencies during inference, balancing between pretrained and extrapolated positional regimes; LongRoPE (Ding et al., 2024): Augments RoPE with specially designed extrapolation functions, enabling robust generalization to long sequences without degrading attention quality. We apply these techniques during fine-tuning, at inference time, or both.
We also incorporate Context Parallelism built upon Ring Attention (Liu et al., 2023a), which cyclically exchanges key-value shards across devices to compute full attention with significantly reduced memory overhead. We apply these techniques during fine-tuning, inference, or both as essential tools for handling long-context inputs and evaluate how effectively they improve practical performance when interleaved trajectories become substantially long.
@@ 5 Experiments
Figure 4: Results of Needle(s) in the Embodied Haystack. The x-axis denotes the input context length (tokens), and the y-axis denotes the needle depth (percentage position of the evidence within the trajectory). The white dashed line denotes the maximum pre-training context length of each model. Qwen2.5-VL was initially pre-trained with an 8K token context window and incrementally scaled up to 32K tokens in subsequent stages (Bai et al., 2025). The gray area indicates contexts not applicable (N/A) due to the model’s smaller image token size which limits sequences to under 128K tokens. Context Parallelism is applied to all experiments with the context sizes over 384K. Full results are provided in Appendix C.1.
Table 3: Architectural Comparison. QA performance on N(s)iEH across different architectures in a setting where the full trajectory context is available. LLaVA-OV model fails when provided with the full trajectory as input (>>1M tokens). GPT-4o is used as an LLM judge for evaluation. For Memory-Augmented approaches, we use a CLIP encoder (Radford et al., 2021) to compute embedding similarity between goal descriptions and historical state observations, selecting the Top-10 and Top-20 most relevant images.
NiEH (Single-evidence) NsiEH (Multi-evidence)
LLaVA-OV Qwen2.5-VL Gemini 2.0 Flash LLaVA-OV Qwen2.5-VL Gemini 2.0 Flash
(a) Full trajectory (Image-only) 0.0 44.67 35.44 0.0 19.34 37.08
(b) Interleaved Goal-State-Action 0.0 51.64 83.04 0.0 43.24 62.72
(c) Memory-Augmented (Text) 51.33 54.38 56.21 53.83 51.33 40.99
(d) Memory-Augmented (Image, Top-10) 35.80 33.99 25.18 22.15 18.63 21.18
(e) Memory-Augmented (Image, Top-20) 37.36 42.68 28.16 23.80 21.46 30.25
@@ 5.1 Static Evaluation: Needle(s) in the Embodied Haystack
Results. Figure 4 (top row) shows the performance of LLaVA-OneVision (OV) 7B across different context extension methods. While Dynamic scaling struggles with contexts beyond 64K tokens, YaRN x4 improves the baseline, handling sequences up to 384K tokens. The second row compares different VLMs: DeepSeek-VL, Qwen2.5-VL, and Gemini 2.0 Flash. Each model processes images into tokens differently–DeepSeek (576 tokens/image), Qwen2.5-VL (121 tokens/image), and Gemini 2.0 Flash (258 tokens/image)–which consequently impacts maximum context lengths when transforming N(s)iEH sequences into tokenized inputs (DeepSeek-VL: 512K, Qwen2.5-VL: 128K, Gemini 2.0 Flash: 256K). DeepSeek-VL fails immediately beyond its 4K pre-training limit. Qwen2.5-VL outperforms DeepSeek but shows clear degradation beyond its 32K pre-training limit. Gemini 2.0 Flash performs robustly up to 32K tokens but performance drops below 40% at 256K, indicating that complex long-range multimodal reasoning remains a challenge even for state-of-the-art models. See Appendix C.1 for more results.
Architectural Comparison. Table 3 presents N(s)iEH task performance across different architectures in a setting where full trajectory context is available. LLaVA-OV fails to process these full trajectories (>>1M tokens), exposing an inherent limitation in its ability to handle long contexts. In contrast, Gemini 2.0 Flash effectively leverages this temporal context. For Memory-Augmented approaches, textual memory demonstrates significantly higher performance than image-based memory, as textual state representations provide denser information within a shorter context size. Notably, we observe that Gemini 2.0 Flash tends to hallucinate object locations in image-only settings when language grounding is not provided (Table 3-(a), (d), and (e)).
Figure 5: Sim2Real Transfer on VSI-Bench. We compare Qwen2.5-VL 7B (Base) against the same model fine-tuned on 226K synthetic NiEH QA pairs. Our model achieves a +11.2% in Object Appearance Order and a +2.1% improvement overall, demonstrating effective transfer to photorealistic environments.
@@
Single vs Multi-evidence Reasoning. Comparing NiEH to the more challenging NsiEH task, we observe a significant performance drop in the multi-evidence setting. This is especially pronounced for questions involving sparse or distant evidence (e.g., “Where was the Mug before you put it on the CounterTop?”) or questions requiring the aggregation of multiple evidences (e.g., “How many times did you move the Apple?”), as shown in Figure 2(b). These results demonstrate that our NiEH and NsiEH tasks pose a substantial challenge to current long-context models and success requires both fine-grained temporal memory and multi-evidence reasoning across extended interactions.
5.2 Online Evaluation: Long-horizon Tasks
We conduct an online evaluation using the AI2THOR simulator. We evaluate models on the final long-horizon task of each test trajectory (e.g., Sub-goal #23 in Figure 1), which requires the agent to utilize history from the preceding hundreds of steps. We fine-tune both LLaVA-OV and Qwen2.5-VL on our training set. Detailed training specifics are available in Appendix B.
Results. Table 4 presents the success rates from our interactive online evaluation. LLaVA-OV failed to complete any long-horizon tasks in this setting (>>1M tokens); therefore, we report results only for the Qwen2.5-VL model in Table 4. We observe that interleaved models consistently outperform Memory-Augmented approaches, suggesting that processing the continuous history is more effective than retrieval-based methods for grounded reasoning. Training with longer contexts leads to significant gains: the 128K context model achieves a 9.2% success rate, confirming that access to extended history is crucial for long-horizon performance. Incorporating long-context extension techniques further boosts performance. YaRN (x4) achieves the highest total success rate of 11.5%, providing a +2.3% gain over the base interleaved model. However, the Put task remains unsolved across all methods, as Put task comes after Go to and Pick up, and the agent rarely reaches this final stage due to failures in the earlier steps. Evaluation results for full-episode execution are available in Appendix C.2. Overall, these results demonstrate that online evaluation for long-horizon tasks remains a significant open challenge.
Table 4: Online Evaluation. Success Rates (%) for the long-horizon tasks. 32K and 128K denote the context size used during fine-tuning.
Method Go to Pick up Put Total
Memory-Augmented (Text) 10.8 1.5 0.0 6.1
Memory-Augmented (Image, Top-20) 6.2 0.0 0.0 3.1
Interleaved Goal-State-Action (32K) 11.3 1.5 0.0 6.4
Interleaved Goal-State-Action (128K) 18.5 2.3 0.0 9.2
+ Dynamic Scaling 20.0 2.3 0.0 9.9
+ YaRN (x4) 23.1 2.6 0.0 11.5
5.3 Demonstrations: Sim2Real Transfer and Low-level Control
Finally, we demonstrate the versatility of $\infty$-Thor through Sim2Real transfer and integration with low-level control policies.
Sim2Real Transfer. We show the transferability of our framework. We used the NiEH pipeline to generate 226K synthetic QA pairs from the training set and fine-tuned the Qwen2.5-VL model. We evaluated the model on VSI-Bench (Yang et al., 2024), a photorealistic benchmark for visual-spatial reasoning. As shown in Figure 5, our model achieves +11.2% improvement on the Object Appearance Order task, confirming that temporal reasoning skills learned in simulation transfer to real-world images. Overall, we observe a +2.1% performance gain with negligible degradation in other spatial tasks.
Low-level Manipulation. We integrated $\infty$-Thor with ManipulaTHOR (Ehsani et al., 2021) to enable fine-grained, low-level robot arm manipulation control. Table 5 reports the success rates on these tasks using OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025). Ego-centric views are provided as image input, and the models generate the next action $(\Delta x,\Delta y,\Delta z)$ to control the robot arm. Overall performance remains low, with both models struggling on manipulation tasks. We attribute these results primarily to two factors: (1) ManipulaTHOR’s egocentric view, which provides nearly identical images for fine-grained arm movements $(\Delta x,\Delta y,\Delta z)$ smaller than 0.01, making it difficult for VLA models to perceive subtle adjustments; and (2) discrepancies between the robot arm model in ManipulaTHOR and those assumed by the VLA models, causing many predicted actions to fail in the physical simulation. We provide more results and analysis in Appendix C.3.
Table 5: Low-level Manipulation Task. Success rates (%) of OpenVLA-7B and SpatialVLA-4B on manipulation tasks in ManipulaTHOR.
Task OpenVLA-7B SpatialVLA-4B
Pick up 16.52 18.70
Put 2.08 1.19
@@ 6 Conclusion
We presented $\infty$-Thor, a new framework for long-horizon embodied tasks designed to advance long-context understanding in embodied AI. Our framework enables scalable synthesis of long, complex trajectories paired with ground-truth action sequences, and supports both offline training and online interaction with the environment. As part of this framework, we introduced a novel embodied QA benchmark, Needle(s) in the Embodied Haystack, that challenges agents to reason over sparse, temporally distant visual evidence embedded within extended trajectories. To equip models for this setting, we explored policy architectures including interleaved Goal-State-Action and Memory-Augmented Goal-State modelings, along with context extension techniques and efficient fine-tuning via Context Parallelism. Our experiments demonstrate that while inference-time extension provides gains, access to longer contexts during training is critical for robust performance. Furthermore, we validated the versatility of $\infty$-Thor through successful Sim2Real transfer to photorealistic benchmarks and integration with low-level manipulation controls. Despite these advances, our results highlight that long-horizon reasoning remains an open challenge, with current state-of-the-art VLMs still struggling to reason over extreme context lengths. We hope our framework encourages further research into the next generation of embodied agents capable of robust long-horizon reasoning and planning.
Impact Statement
This paper introduces a framework for training and evaluating long-horizon embodied agents, with novel benchmarks, architectural explorations, and training methodologies. We view our framework as a platform to teach embodied agents how to reason over long temporal horizons and execute complex, multi-step instructions in dynamic environments.
The most relevant downstream task for this work is long-horizon embodied planning and robotic manipulation. As our dataset is synthesized within the AI2-THOR simulator, there is a risk of domain gap and limited diversity in scene configurations, which may result in performance biases when applied to unstructured real-world environments.
@@ A.2 Constructing Long-Horizon Trajectories
Figure 6: Example of the trajectory and a long-horizon embodied task generated from $\infty$-Thor. The final goal (“Put the tomato on the counter top” at t=670) requires recalling both the tomato (seen at t=17) and the counter (seen at t=560) to solved the long-horizon task. Context size refers to the input token length when converting the trajectory into the VLM input space.
@@
13:  {Final sub-task with long-term object dependency}
@@
Figure 6 shows an example of the trajectory and a long-horizon embodied task generated from $\infty$-Thor. To synthesize long-horizon trajectories, we construct each trajectory by sequentially chaining successful sub-tasks sampled from a predefined set of task templates. This process is illustrated in Algorithm 1. We begin by sampling a task template from a fixed task pool, which includes goal types such as pick and place simple, pick two obj and place, and pick and place with movable recep. Each sampled template requires relevant objects in the scene (e.g., pickupable items, target receptacles), which are then used to define the goal for that task.
A.3 Dataset Statistics
Table 8 summarizes the statistics for the NiEH test set and the generated long-horizon trajectories. The training set consists of 2,456 trajectories with an average of 14 sub-goals and 405 steps per episode. Token lengths are calculated using the LLaVA-OneVision tokenizer. With an average input length of 602K tokens, the entire training dataset amounts to approximately 1.48B tokens (2,456 $\times$ 602K), providing a substantial corpus for learning long-horizon temporal dependencies.
Table 8: Dataset statistics
. NiEH testset Single-clue Multi-clue # of question-answer pair 829 474 Trajectory Train Dev Test # trajectory 2,456 125 225 # avg subgoals 14 16 18 # max subgoals 30 24 33 # avg steps 405 613 627 # max steps 654 890 952 # avg token length 602K 880K 912K # max token length 954K 1.2M 1.3M # avg interaction with objects per episode 31 36 37 # max interaction with objects per episode 82 62 74 # scene 82 30 39
Appendix B Training Details
We fine-tune the LLaVA-OneVision 7B and Qwen2.5-VL 7B model on our training set while freezing the vision encoder. Training utilized 8 H100 GPUs with tensor (Shoeybi et al., 2020) and pipeline parallelism (Huang et al., 2019) for 32K context, and Context Parallelism for larger contexts (64K, 130K). The model is trained using a next-action prediction objective, where only the action tokens are optimized, conditioned on the goal and state tokens. Table 9 summarizes the training specifications for different context lengths. For 32K training, we apply tensor parallelism with a degree of 4 and pipeline parallelism with a degree of 2, utilizing 8 H100 GPUs in total. Since pipeline parallelism requires the batch size to match the pipeline degree, we set the batch size to 2. For longer context lengths, we use context parallelism: 8-way for 64K (on 8 GPUs) and 16-way for 130K (on 16 GPUs). All models are fine-tuned for approximately 3 epochs with a learning rate of 1e-5, using the AdamW optimizer and a linear learning rate schedule with a 0.03 warmup ratio.
@@ Appendix B Training Details
Appendix C Additional Results
C.1 Static Evaluation: Needle(s) in the Emboided Haystack
Figure 7: Full results of Needle(s) in the Embodied Haystack. The x-axis denotes the input context length (tokens), and the y-axis denotes the needle depth. The white dashed line denotes the maximum pre-training context length of each model.
Figure 8: Agent’s reward across different experimental configurations for high-level planning tasks. We compare (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, and (c) fine-tuning with different context lengths using Context Parallelism. (d) summarizes the most effective strategies, highlighting that exposure to longer contexts during training significantly improves performance. Non-planner models cannot generate valid actions after around 250 steps ($\approx$376K in context size).
Figure 7 presents the performance of LLaVA-OneVision (OV) 7B (Li et al., 2024a) model across various context extension methods. Linear Interpolation, Dynamic Scaling, and LongRoPE scaling all struggled with very long contexts beyond 128K tokens (the results of Linear Interpolation are excluded from Figure since it fails at all examples). YaRN consistently outperformed other methods, effectively handling contexts above 384K tokens, likely due to its architectural alignment with LLaVA-OV’s Qwen2 LM backbone, which employs RoPE and YaRN scaling during pretraining. YaRN performed best at moderate scaling factors (e.g., x4), however, further scaling to x8 and x16 did not yield additional gains. In particular, x16 slightly improved performance in the 256K–384K token range but led to degradation notably at shorter context sizes (<<64K), suggesting that excessive scaling may introduce instability and negatively impact performance. Overall, all methods fail beyond 512K tokens, highlighting a need for improved long-context methods. The third row includes additional results from DeepSeek-VL 7B (Lu et al., 2024), Qwen2.5-VL 7B (Bai et al., 2025), and Gemini 2.0 Flash [^1] We used the gemini-2.0-flash-001 version for all experiments.. Although Gemini is renowned for strong long-context multimodal reasoning (up to 3,000 image input), our results (<<40% with full trajectories at 256K) indicate substantial room for improvement.
C.2 Online Evaluation: High-level Planning
Plan-level Evaluation. We evaluate agent performance using a plan-level framework, where each plan corresponds to a short sequence of actions aimed at achieving a specific intermediate sub-goal (e.g., navigating to an object, placing an item). A trajectory is composed of multiple such plans, executed sequentially. For the interactive evaluation, the agent is presented with the current plan’s goal along with the history of previous GT states and actions. Using this context, the agent predicts the next action and interacts step-by-step with the environment. The interaction continues until the current plan is either successfully completed or terminated due to failure (e.g., collisions or deadlocks). After each plan, the context is reset to include the GT actions and states from the completed portion of the trajectory, and the agent proceeds to the next plan. This ensures that each plan is evaluated independently, conditioned only on the correct prior history. The agent’s performance is measured via cumulative reward across all plans in the trajectory. Pseudocode for this evaluation procedure is provided in Algorithm 2.